# Grounding Fallacies Misrepresenting Scientific Publications in Evidence

### v1

@@ 7.1 Fine-Tuned Models for Scientific AFC
Given a claim $\overline{c}$ that misrepresents the publication $S$, we form $n$ fact-checking instances ($\overline{c}$, $S_{j}$) by treating each of the $n$ annotated passages $S_{j}\in S$, as provided evidence. We use the scientific AFC models used from §5 to produce $n$ veracity labels given $\overline{c}$ and each of the selected passages $S_{j}$. Following Schlichtkrull et al. (2023), we assign an overall veracity label for a claim as Mixed if the labels Supported and Refuted were predicted. If the model only predicted NotEnoughInformation (NEI), the overall verdict was NEI. In all other cases in which the model predicted a mix of Supported (with optional NEI) or Refuted (with optional NEI) we label it as Supported or Refuted respectively. We note that the studies in MissciPlus constitute trustworthy evidence. They do not support the inaccurate claims that HFC has rated as misrepresenting these studies. A literate scientific AFC model should equally detect the claims as misinformation based on the same trustworthy evidence.
@@
Table 5 shows that all AFC models yielded high in-domain accuracy on the respective AFC dataset (AFC Acc.) ranging from 82-90 points in accuracy. Because Missci assumes that the study provides sufficient evidence for the claim, any prediction other than Supported can be considered as correctly rejecting the claim. Yet, AFC systems falsely predict 37-56% of the claims to be true. The seemingly best performing CovidFact model uses binary classification without NEI and benefits from a reduced chance to predict the overall label Supported as per the overall label assignment.
The evaluation on misinformation favors sceptical models regardless of whether a passage truly undermined a claim. Grounding the verdict of AFC models is critical for the trustworthiness and often part of the evaluation protocol (Thorne et al., 2018; Wadden et al., 2020; Glockner et al., 2024b). To understand whether the AFC models reject the claims based on the undermining evidence that indicates reasoning gaps, we visualize the AFC prediction over different passages in Figure 5. Intuitively, the models mostly predict Supported over passages based on which the claim was made (top left). If the same passage is additionally linked to a reasoning gap (bottom left), the predicted distribution does not change much, suggesting unawareness of the mismatches. Based on passages that are linked to a reasoning gap but not to the accurate premise (top right), the distribution changes towards NEI. This is similar to completely unrelated passages (bottom right) and suggests that AFC models struggle to detect scientific distortions as such. We provide additional analysis in §H.1.
@@
GPT3.5 1.2 71.0 28.2 46.7 17.3 35.7
Llama2 0.0 100.0 0.0 29.7 69.3 3.0
Know Llama3 8.3 88.9 2.4 68.7 26.0 3.7
(Ask) GPT4 3.6 68.3 27.8 49.7 36.3 6.7
@@
7.2 LLMs as Scientific AFC Models
Experimental Setup
We tested how well LLMs can fact-check claims in MissciPlus when prompted to predict their veracity directly. We prompted each LLM (a) without any evidence to test whether it already has parametric knowledge about the claims from pretraining (Magar and Schwartz, 2022) and (b) with the relevant evidence passages to test their reasoning capabilities in a RAG style, similarly to Manakul et al. (2023). For parametric knowledge, we let the LLM generate a fact-checking article (FC), which is closest to our application, but may bias the model due to the prevalence of misinformation in fact-checking articles. As a neutral prompt, we additionally asked the LLM to predict the veracity to the best of their knowledge. All prompts are listed in §I.
Results
Table 6 shows the results on MissciPlus and 100 randomly sampled correct claims from HealthVer and CovidFact (50 from each dataset), both of which focus on real-world health-related claims from the internet. Note that the percentages do not always add up to 100% as LLMs can refuse to answer. The results suggest that all LLMs seem to have learned that the claims in MissciPlus are not true. Also, true claims are rarely classified as False and mostly as True. The proportion of claims predicted as True increases when providing additional evidence. Interestingly, this also happens to misinformation in MissciPlus despite the LLM seemingly knowing that these claims are not true and despite that the misrepresented evidence does not truly support the claims. This is similar to overruling the internal knowledge with evidence (Wu et al., 2024) and suggests that the LLMs do not identify the reasoning gaps between the claim and the evidence, as outlined by the HFCs. Our findings indicate that LLMs, similarly to humans, are prone to misinterpreting scientific publications. Due to their high persuasiveness (Augenstein et al., 2023; El-Sayed et al., 2024), this can lead to disastrous problems when people rely on LLMs, even if LLMs transparently output the (misrepresented) used evidence.
Discussion
Similarly to previous work on fallacy recognition (Jin et al., 2022; Alhindi et al., 2022; Glockner et al., 2024a), the prompts in this work focus on which fallacies apply instead of whether any applies, and are no panacea either: In preliminary experiments §H.2, we prompt LLMs to verbalize the fallacies or output that there is “no fallacy”. The LLMs found fallacies in 85-99% of the claims in MissciPlus but also in 78-99% of the true claims. MissciPlus unifies fallacy detection with AFC and provides a resource for studying how to handle fallacious arguments and true claims alike.
@@ 8 Conclusion and Future Work
We introduced MissciPlus, a new dataset for reconstructing fallacious arguments based on realistic evidence. We showed that existing ranking models and LLMs struggle to reconstruct fallacious arguments using the evidence that reveals reasoning gaps. Moreover, fine-tuned AFC models and LLMs failed to refute claims in MissciPlus when presented with misrepresented evidence, highlighting the dangers of persuasive LLMs.
In future work we plan to leverage MissciPlus to build and evaluate models that are resistant to the nuanced mismatches between claims and their corresponding cited evidence documents.
@@
MissciPlus assumes that the claims are based on a single publication only and that each publication is inherently trustworthy, i.e., that the only error was done in the reasoning from the publication to the claim. In the real world, finding complementary credible evidence is critical. We found that single passages are insufficient to ground each fallacy in evidence and ignore content that is multi-model, requires multiple passages or can only be found from the entire study. Our evaluation is based on automatic matching, which is inevitably imperfect. Models may detect fallacies that are valid but not contained in the original annotations in Missci. We compensate for this with two complementary matching strategies and experiments over different seeds and prompt variations to confirm the robustness of our observations. By focusing on recall, we further do not penalize models for prediction fallacies that are not within our annotation while still requiring the models to detect the most prominent ones as highlighted by the annotators in Missci (from the HFC articles).
@@
False interpretations of health-related claims can have disastrous effect. Any output of models derived from MissciPlus only serves research purposes to detect such misinformation, but under no circumstances must be considered accurate without consulting experts in the field. Our work poses dangers for dual use, particularly verbalizing the fallacious reasoning to draw incorrect conclusions from real-world studies. While generating (parts of) misinformation always poses a risk, it is unavoidable to build resilience against real-world misinformation, as demonstrated in previous work (Zellers et al., 2019; Huang et al., 2023; Alhindi et al., 2023; Glockner et al., 2024a).
@@
All publications used in MissciPlus have been published by the respective authors and we did not anonymize their work. All publications used in MissciPlus are openly available and are part of the public discourse; in fact, they have even been distorted by misinformation. Hence, similarly to other scientific corpora (Lehman et al., 2019; Lo et al., 2020) that rely on such publications as evidence, we did not ask for explicit permission from the authors of each study to use their work..
@@ Acknowledgments
This research work has been funded by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE and by the German Research Foundation (DFG) as part of the UKP-SQuARE project (grant GU 798/29-1). Yufang Hou is supported by the Visiting Female Professor Programme from TU Darmstadt. We gratefully acknowledge the support of Microsoft with a grant for access to OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research). We are grateful to our dedicated annotators who helped to create MissciPlus. Finally, we wish to thank Jan Buchmann, Sukannya Purkayastha, Luke Bates and Jonathan Tonglet for their valuable feedback on an early draft of this work.
@@ Appendix A Fallacies in MissciPlus
Appendix B Dataset Construction
B.1 Passage Extraction
@@ B.3 Annotation
B.4 Missing Passage Link Analysis
@@ B.5 Licenses
Appendix C Length of Passages in MissciPlus
@@
Appendix D Scientific AFC Models
@@ D.2 Hyperparameter search
Appendix E Details on Finding the Kernel of Truth
E.1 Implementation Details
E.1.1 Baselines
As baselines, we randomly shuffle all passages (random) using five different seeds (1-5) or order passages as they appear within the original publication (ordered). We rank passages via BM25888https://pypi.org/project/rank-bm25/ for ranking based on lexical similarity. For pre-processing, we use SpaCy [^9] https://spacy.io/ for tokenization and converting all tokens to lowercase.
E.1.2 Embedding-based approaches
@@
We experiment with various sentence-embedding models (in addition to those reported in Table 1). This includes SBERT101010all-mpnet-base-v2 (Reimers and Gurevych, 2019), and sentence transformers fine-tuned on BioBERT (Lee et al., 2020), PubMedBERT (Gu et al., 2021), and SapBERT (Liu et al., 2021) as provided by Deka et al. (2022). For the INSTRUCTOR, we use the following prompts for the INSTRUCTOR (Su et al., 2022), which follows the official templates:
@@
For all embedding models, we ranked the passages according to the cosine similarity to the embedded claim. For SPICED-IMS, we use the model [^11] https://github.com/copenlu/scientific-information-change provided by Wright et al. (2022) and rank passages according to their IMS between claim and passage. We treat the scientific AFC models from §D as cross-encoder models to jointly encode each claim-passage pair and re-rank passages based on the predicted probability for the label Supported.
@@ E.1.3 LLMs with PRP
For PRP, we use the claim and two passages as input to the LLM and prompt it to output which passage should be ranked first. Following Qin et al. (2023), we use these outputs to re-rank passages akin to bubble sort. The bubble sort algorithm ensures that the top $k$ elements are ranked at the top after $k$ iterations. We evaluate four prompts on the validation split in Table 17, starting with the prompt provided by Qin et al. (2023) and three deviations tailored towards our task. We further assess whether passages should include or exclude the title and found that including the title generally improves performance. During prompt selection in Table 17 we randomly shuffle all passages prior to reranking them to avoid the impact of the strong positional bias on the validation split (P@$1$: 0.577). Since our metrics are only sensitive to the top-ranked results, PRP can only be completed over a few iterations to reduce costs without affecting the performance much. Following our experiments on the validation split, we only ran PRP for three iterations on the test set. Prompts are listed in §I.1.
E.2 Exhaustive Results on MissciPlus
@@
Appendix F Details on Finding Undermining Passages
We only evaluate finding undermining passages in the open subset. Hence, the results represent a lower bound of the system’s performance. Per our hyper-parameter search (§F.1.1 and §F.1.2), we always provide one randomly sampled passage $S^{0}$ to the model, which is necessary to understand the rationale (and reasoning gaps) behind the claim. If no passage $S^{0}$ exists, we use the paraphrased accurate premise from Missci instead. All experiments are averaged over five seeds (1-5).
F.1 Implementation Details
F.1.1 AFC as Rankers
@@
F.1.2 Dense embedding rankers
@@
Using the same ranking models via embeddings or BM25 as in §E, we assess the impact of prepending a passage $S^{0}$ to the claim on the validation split in Table 20. We select the best configuration according to mAP for the test split. For the INSTRUCTOR, we modify the prompt to encode the claim by looking for “refuting” instead of “supporting” sentences (cf. §E.1.2), which resulted in a higher MAP on the validation set.
F.2 Impact of $S^{0}_{j}$ Passages
@@
F.3 Exhaustive Results on MissciPlus

### v2

@@ 7.1 Fine-Tuned Models for Scientific AFC
Given a claim $\overline{c}$ that misrepresents the publication $S$, we form $n$ fact-checking instances ($\overline{c}$, $S_{j}$) by treating each of the $n$ annotated passages $S_{j}\in S$, as evidence. We use the scientific AFC models from §5 to predict $n$ veracity labels. Following Schlichtkrull et al. (2023), we assign an overall veracity label for a claim as Mixed if the labels Supported and Refuted were predicted. If the model only predicted NotEnoughInformation (NEI), the overall verdict was NEI. We label all other cases in which the model predicted Supported (with optional NEI) or Refuted (with optional NEI) as Supported or Refuted, respectively. The studies in MissciPlus constitute trustworthy evidence that do not support the inaccurate claims – HFC have rated them to misrepresenting these studies. A literate scientific AFC model should equally detect the claims as misinformation.
Table 5 shows that all AFC models yielded high in-domain accuracy (82-90%) on the respective AFC dataset. In MissciPlus, the misrepresented studies were (mis)used as evidence to back up inaccurate claims. Therefore, if the AFC model assigns any label other than Supported, it can be considered a correct rejection of the claim. Yet, AFC systems falsely predict 37-56% of the claims to be true. The seemingly best CovidFact model uses binary classification without NEI and benefits from a substantially increased chance to predict Mixed.
Grounding the verdict of AFC models in the used evidence is critical for their trustworthiness and often is a key part of the evaluation protocol (Thorne et al., 2018; Wadden et al., 2020; Glockner et al., 2024b). Table 5 only shows if the LLM rejected a claim based on any passage, not if the LLM assigned the correct label for the provided passage. To understand whether the AFC models reject the claims based on the undermining evidence that indicates reasoning gaps, we visualize the AFC prediction over different passages in Figure 5. Intuitively, the models mostly predict Supported over passages based on which the claim was made (top left). If the same passage is additionally linked to a fallacy (bottom left), the predicted distribution does not change much, suggesting unawareness of the fallacy. Based on passages that are linked to a fallacy but not to the accurate premise (top right), the distribution changes towards NEI. This is similar to completely unrelated passages (bottom right), again suggesting unawareness of the scientific distortion. More analysis is in §I.1.
@@
7.2 LLMs as Scientific AFC Models
HFCs disseminates their fact-checking articles widely, which can give unfair advantages to models when they are evaluated on the same misinformation (Glockner et al., 2022). We prompted each LLM (a) without any evidence to test whether it already has parametric knowledge about the claims from pretraining (Magar and Schwartz, 2022) and (b) with the relevant evidence passages to test their reasoning capabilities in a RAG style, similarly to Manakul et al. (2023). For parametric knowledge, we let the LLM generate a fact-checking article (FC), which is closest to our application, but may bias the model to refute a claim due to the prevalence of misinformation in fact-checking articles. As a neutral prompt, we additionally asked the LLM to predict the veracity of the claim to the best of their knowledge. All prompts are listed in §J. Table 6 shows the results on MissciPlus and 100 randomly sampled correct claims from HealthVer and CovidFact (50 per dataset). Both datasets contain real-world health-related claims from the internet. The percentages may not add up to 100% if LLMs refuse to answer.
@@ 7.2 LLMs as Scientific AFC Models
GPT3.5 0.8 71.0 28.2 46.7 17.3 35.7
Llama2 0.0 100.0 0.0 29.7 69.3 1.0
Know Llama3 8.3 88.9 2.4 68.7 26.0 3.0
(Ask) GPT4 3.6 68.3 27.8 49.7 6.7 36.3
@@
All LLMs tend to have parametric knowledge about the veracity of many claims. When providing evidence, the ratio of claims predicted as True increases not only for correct claims but also for misinformation, despite that the misrepresented evidence does not truly support the claims. This phenomenon is similar to overruling the internal knowledge with evidence (Wu et al., 2024) and suggests that the LLMs do not identify the reasoning gaps between the claim and the evidence. Similarly to humans, LLMs seem prone to misinterpreting scientific publications. Due to their high persuasiveness (Augenstein et al., 2024; El-Sayed et al., 2024), this can lead to disastrous problems when people rely on LLMs, even if LLMs transparently output the (misrepresented) used evidence.
Akin to previous work on fallacy recognition (Jin et al., 2022; Alhindi et al., 2022; Glockner et al., 2024a), the prompts in this work focus on which fallacies apply instead of whether any applies, and are no panacea either: in preliminary experiments (cf. §I.2), LLMs found fallacies in 85-99% of the claims in MissciPlus but also in 78-99% of the true claims. Overall, MissciPlus unifies fallacy detection with AFC and provides a resource for studying how to handle fallacious arguments and true claims alike.
@@ 8 Conclusion and Future Work
We introduced MissciPlus, an extension of Missci to reconstruct fallacious arguments based on real passages of the misrepresented studies. We showed that existing ranking models and LLMs struggle to reconstruct fallacious arguments using the real-world evidence. Moreover, fine-tuned AFC models and LLMs failed to refute claims in MissciPlus when presented with misrepresented evidence, highlighting the dangers of persuasive LLMs. Future work may improve models using synthetic data generation approaches, or extend the task definition over multiple documents.
@@
MissciPlus assumes that the claims are based on a single publication only and that each publication is inherently trustworthy, i.e., that the only error was done in the reasoning from the publication to the claim. In the real world, finding complementary credible evidence is critical. We observed that sometimes single passages are insufficient to ground each fallacy in evidence. Currently MissciPlus does not address grounding fallacies in multi-modal content, cases that require evidence from multiple passages, or conclusions that can only be drawn from the analysis of the entire study. We leave these challenges for future research. Our evaluation is based on automatic matching, which is inevitably imperfect. Models may detect fallacies that are valid but not contained in the original annotations in Missci. We compensate for this with two complementary matching strategies and experiments over different seeds and prompt variations to confirm the robustness of our observations. By focusing on recall, we further do not penalize models for predicting fallacies that are not within our annotations while still requiring the models to detect the most prominent ones as highlighted by the annotators based on the HFC articles in Missci. Our results are reported over two representative state-of-the-art LLMs at the time of writing (Llama3 8B as an easy-to-run open-source LLM and GPT4 as a proprietary LLM), and our claims are bound to these models. We opted for extensive tuning of these models to establish strong baselines for the novel tasks rather than providing comprehensive comparisons across various LLMs. While our focus lies in grounding the individual argument constituents in the real-world misrepresented study, which contains 2,257 annotated links and 400 argument constituents, MissciPlus is only based on 114 fallacious arguments. This may lead to variance and biases in the experiments, which must be interpreted with caution. We note that creating high-quality fallacy datasets with complex fallacious arguments requires suitable, professionally fact-checked claims, for which data is scarce. Future work could explore synthetic data generation to help bridge this gap.
@@
False interpretations of health-related claims can have disastrous effect. Any output of models derived from MissciPlus only serves research purposes to detect such misinformation, but under no circumstances must be considered accurate without consulting experts in the field. Our work poses dangers for dual use, particularly verbalizing the fallacious reasoning to draw incorrect conclusions from real-world studies. While generating (parts of) misinformation always poses a risk, it is unavoidable to build resilience against real-world misinformation, as demonstrated in previous work (Zellers et al., 2019; Huang et al., 2023; Alhindi et al., 2024; Glockner et al., 2024a).
@@
All publications used in MissciPlus have been published by the respective authors and we did not anonymize their work. All publications used in MissciPlus are openly available and are part of the public discourse; in fact, they have even been distorted by misinformation. Hence, similarly to other scientific corpora (Lehman et al., 2019; Lo et al., 2020) that rely on such publications as evidence, we did not ask for explicit permission from the authors of each study to use their work.
@@ Acknowledgments
This research work has been funded by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research, Science and the Arts within their joint support of the National Research Center for Applied Cybersecurity ATHENE. Yufang Hou was supported by the Visiting Female Professor Programme from TU Darmstadt. We gratefully acknowledge the support of Microsoft with a grant for access to OpenAI GPT models via the Azure cloud (Accelerate Foundation Model Academic Research). We are grateful to our dedicated annotators who helped to create MissciPlus and to the anonymous reviewers and meta reviewer for their valuable feedback. Finally, we wish to thank Jan Buchmann, Sukannya Purkayastha, Luke Bates and Jonathan Tonglet for their valuable feedback on an early draft of this work.
@@ Appendix A Fallacies in MissciPlus
Appendix B Example Argument
Figure 6: A complete fallacious argument, including three (out of 29) passages from the misrepresented study. The arrows indicate which original passage in MissciPlus corresponds to the paraphrased publication context in Missci. Some passages convey the same content in different publication contexts, and the same content may appear in multiple passages.
We present a complete fallacious argument, including three (out of 29) original passages (blue) from the misrepresented study, in Figure 6. The claim employs three distinct fallacies (red), each of which can be identified based on different information from the study. Two of the displayed passages contain relevant information for detecting these fallacies. The middle passage provides all the necessary information to identify all three fallacies, as well as the accurate premise. The first passage enables detection of the fallacies related to the study’s investigation being limited to cell cultures ($s_{1}$) and its use of chloroquine rather than hydroxychloroquine ($s_{3}$). In MissciPlus, only the original passages (marked in blue) are provided to the model, which must infer the fallacies from the complex scientific content. Although not used in this work, MissciPlus includes sentence-level annotations.
Appendix C Dataset Construction
C.1 Passage Extraction
@@ C.3 Annotation
C.4 Missing Passage Link Analysis
@@ C.5 Licenses
Appendix D Length of Passages in MissciPlus
@@
Appendix E Scientific AFC Models
@@ E.2 Hyperparameter search
Appendix F Details on Finding the Kernel of Truth
F.1 Implementation Details
F.1.1 Baselines
As baselines, we randomly shuffle all passages (random) using five different seeds (1-5) or order passages as they appear within the original publication (ordered). We rank passages via BM25666https://pypi.org/project/rank-bm25/ for ranking based on lexical similarity. For pre-processing, we use SpaCy [^7] https://spacy.io/ for tokenization and converting all tokens to lowercase.
F.1.2 Embedding-based approaches
@@
We experiment with various sentence-embedding models (in addition to those reported in Table 1). This includes SBERT888all-mpnet-base-v2 (Reimers and Gurevych, 2019), and sentence transformers fine-tuned on BioBERT (Lee et al., 2020), PubMedBERT (Gu et al., 2021), and SapBERT (Liu et al., 2021) as provided by Deka et al. (2022). For the INSTRUCTOR, we use the following prompts for the INSTRUCTOR (Su et al., 2023), which follows the official templates:
@@
For all embedding models, we ranked the passages according to the cosine similarity to the embedded claim. For SPICED-IMS, we use the model [^9] https://github.com/copenlu/scientific-information-change provided by Wright et al. (2022) and rank passages according to their IMS between claim and passage. We treat the scientific AFC models from §E as cross-encoder models to jointly encode each claim-passage pair and re-rank passages based on the predicted probability for the label Supported.
@@ F.1.3 LLMs with PRP
For PRP, we use the claim and two passages as input to the LLM and prompt it to output which passage should be ranked first. Following Qin et al. (2024), we use these outputs to re-rank passages akin to bubble sort. The bubble sort algorithm ensures that the top $k$ elements are ranked at the top after $k$ iterations. We evaluate four prompts on the validation split in Table 17, starting with the prompt provided by Qin et al. (2024) and three deviations tailored towards our task. We further assess whether passages should include or exclude the title and found that including the title generally improves performance. During prompt selection in Table 17 we randomly shuffle all passages prior to reranking them to avoid the impact of the strong positional bias on the validation split (P@$1$: 0.577). Since our metrics are only sensitive to the top-ranked results, PRP can only be completed over a few iterations to reduce costs without affecting the performance much. Following our experiments on the validation split, we only ran PRP for three iterations on the test set. Prompts are listed in §J.1.
F.2 Exhaustive Results on MissciPlus
@@
Appendix G Details on Finding Undermining Passages
We only evaluate finding undermining passages in the open subset. Hence, the results represent a lower bound of the system’s performance. Per our hyper-parameter search (§G.1.1 and §G.1.2), we always provide one randomly sampled passage $S^{0}$ to the model, which is necessary to understand the rationale (and reasoning gaps) behind the claim. If no passage $S^{0}$ exists, we use the paraphrased accurate premise from Missci instead. All experiments are averaged over five seeds (1-5).
G.1 Implementation Details
G.1.1 AFC as Rankers
@@
G.1.2 Dense embedding rankers
@@
Using the same ranking models via embeddings or BM25 as in §F, we assess the impact of prepending a passage $S^{0}$ to the claim on the validation split in Table 20. We select the best configuration according to mAP for the test split. For the INSTRUCTOR, we modify the prompt to encode the claim by looking for “refuting” instead of “supporting” sentences (cf. §F.1.2), which resulted in a higher MAP on the validation set.
G.2 Impact of $S^{0}_{j}$ Passages
@@
G.3 Exhaustive Results on MissciPlus